please help me to solve this problem by completing the sqaure step be step x^2+10x-3=0
Answer by ewatrrr(24785)


Hi
completing the square
x^2+10x-3=0
x^2+10x -3 = 0
(x+5)^2 -25 -3 = 0 |Note:(x+5)^2 = 'x^2 + 10x' + 25 (Thus the -25)
(x+5)^2 -28 = 0